OpenStudy (anonymous):

What is the product of the two smallest prime factors of 2^1024-1?

OpenStudy (amistre64):

1*2?

OpenStudy (anonymous):

1 is not prime.

OpenStudy (amistre64):

its prime enough :)

OpenStudy (amistre64):

ohno! a wikipedia proof ...

OpenStudy (amistre64):

The number 1 is a special case which is considered neither prime nor composite (Wells 1986, p. 31). Although the number 1 used to be considered a prime (Goldbach 1742; Lehmer 1909, 1914; Hardy and Wright 1979, p. 11; Gardner 1984, pp. 86-87; Sloane and Plouffe 1995, p. 33; Hardy 1999, p. 46), it requires special treatment in so many definitions and applications involving primes greater than or equal to 2 that it is usually placed into a class of its own. - from someplace other than my head i hope

OpenStudy (anonymous):

Also: 2 is clearly not a prime factor of \[ 2^{1024} - 1 \]

OpenStudy (amistre64):

back in my day 1 was a prime ;)

OpenStudy (turingtest):

\[2^{1024}-1=(2^{512}-1)(2^{512}+1)\]i think is a nice direction to go it continues that way

OpenStudy (amistre64):

i just simplified to: 2^1023

OpenStudy (turingtest):

but keep doing difference of squares...

OpenStudy (turingtest):

it has to a reasonable factor evenutally

OpenStudy (turingtest):

Fool has some brilliant solution I klnow

OpenStudy (anonymous):

According to this: http://www.alpertron.com.ar/ECM.HTM \[2^{1024} - 1 = 3 \times 5 \times 17 \times 257 \times 641 \times 65537 \times 274177 \times 2 424833 \times 6 700417 \]\[\times 67 280421 310721 \times 1238 926361 552897 \times 59649 589127 497217 \]\[ \times 5704 689200 685129 054721 \]\[ \times 7 455602 825647 884208 337395 736200 454918 783366 342657 \]\[\times 93 461639 715357 977769 163558 199606 896584 051237 541638 188580 280321 \]\[\times 741 640062 627530 801524 787141 901937 474059 940781 097519 023905 821316 144415 759504 705008 092818 711693 940737 \]... So it would be 3 times 5 = 15. I dont really know how to do it by hand though.

OpenStudy (anonymous):

ya, its 15.

OpenStudy (anonymous):

You should probably just somehow prove that 3 and 5 are prime factors of this number. I do not see a clever method to factorize it by hand since it is so gigantic.

OpenStudy (turingtest):

it factors by difference of squares down to\[(2^2-1)(2^2+1)\cdot\]is how you do it!

OpenStudy (turingtest):

times a bunch of other stuff...

OpenStudy (anonymous):

My apologies my answer was wrong :(

OpenStudy (turingtest):

keep dividing the exponent by 2 in every difference of squares factoring and you will get my answer\[(2^2-1)(2^2+1)(2^4+1)(2^8+1)\cdots\]

OpenStudy (anonymous):

lol, 2^1024-1= (2^512 +1)(2^256 +1)(2^128 +1)(2^64 +1)(2^32+1)(2^16+1) (2^8+1)(2^4 +1)(2^2 +1)(2-1)(2+1)

OpenStudy (turingtest):

yep :D I used to suck at these things, but with you guys around I guess I'm getting better!

OpenStudy (anonymous):

we only need (2^2 +1)(2-1)(2+1) = 15

OpenStudy (anonymous):

Turing I am having a bad day : http://openstudy.com/users/foolformath#/updates/4f367ecde4b0fc0c1a0ce793

OpenStudy (turingtest):

I wish that was how my bad days were I can't even add on mine!

OpenStudy (anonymous):

haha I don't believe you Turing

OpenStudy (anonymous):

Turing here is another way for solving the same problem http://openstudy.com/study#/updates/4f376f36e4b0fc0c1a0d4f7c